OpenStudy (anonymous):

for my homework it says use symbolic integration utiltiy to find the indefinite integral... whats that mean?

OpenStudy (anonymous):

The term symbolic is used to distinguish this problem from that of numerical integration, where the value of F at a particular input or set of inputs, rather than a general formula for F, is sought.

OpenStudy (anonymous):

in general , it just mean "indefinite integral"

OpenStudy (anonymous):

do you have an example of the problem?

OpenStudy (anonymous):

find the indefinite integral of x^3/(1-x^4)^(1/2)

OpenStudy (anonymous):

\[ \int\limits_{}{x} ^{3}/\sqrt{1-x^{4}} dx\]

OpenStudy (anonymous):

substitution: u=1-x^4 then du=-4*x^3 dx try it

OpenStudy (anonymous):

whats that suppose to be?

OpenStudy (anonymous):

x^3 dx =1/4 du right?

OpenStudy (anonymous):

i did it

OpenStudy (anonymous):

i got -1/2(1-x^4)^1/2 +C

OpenStudy (anonymous):

sorry i didnt respont i was looking throug my notes

OpenStudy (anonymous):

\[=-1/4 \int\limits_{ }^{}du/u ^{1/2}=-1/4 * 2/3 * u ^{3/2}=\]

OpenStudy (anonymous):

yes i did that

OpenStudy (anonymous):

=-1/6 (1-x^4) ^3/2 + const

OpenStudy (anonymous):

its not 3/2 because its -1/2

OpenStudy (anonymous):

oups... you are right!

OpenStudy (anonymous):

the square root was originally to the bottom sorry if i didnt make it clear

OpenStudy (anonymous):

thank you ! your help is greatly appreciated :)

OpenStudy (anonymous):

so I got: -1/2 (1-x^4)^1/2 + const right?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

okay lol